# A highly accurate procedure for computing globally optimal Wannier functions in one-dimensional crystalline insulators

### v3

@@ 4 Constructing optimally localized Wannier functions
The construction can be divided into three stages. First, we compute the Fourier series of $V$ as in (53) and transform (32) into its Fourier version in (56). Then, we compute the energy $E_{0}$ and the eigenvector $\mathbf{y}_{0}$ at $k=-\Omega/2$ for the selected band. Second, we solve the ODE defined by (60) and (61) in $[-\Omega/2,\Omega/2]$ with initial conditions given by $E_{0}$ and $\mathbf{y}_{0}$ at $k=-\Omega/2$. From $\mathbf{y}(-\Omega/2)$ and $\mathbf{y}(\Omega/2)$, we determine a constant $\phi_{0}$ to enforce (100) and the Zak phase $\phi_{\rm zak}$ in (68). This is followed by carrying out a gauge transform to obtain the new $\widetilde{\mathbf{y}}$ in (76) with the Berry phase $\phi$ given by (80). Using the definition of $\widetilde{\mathbf{y}}$ in (77), we obtain a function $\widetilde{\alpha}$ in $\mathbb{R}$ that defines the Fourier transform of the Wannier function by (49). Finally, we Fourier transform $\widetilde{\alpha}$ to obtain the Wannier function $W_{0}$ by (48). This Wannier function is guaranteed to be optimal by Theorem 3.12 and Corollary 3.14, and real by Theorem 3.16.
@@
Remark 4.1.
There are various schemes resembling the one above. For example, [20] and Section IV.C in [15] describe a conceptually similar algorithm. Instead of solving the ODE in the second stage above, $u_{k}$, for varying values of $k$ on a grid in $I_{\mathrm{bz}}$, are computed independently and an approximate SVD-based approach is used to fix the broken phase. For a large number of grid points, the SVD-based approach should yield a decent approximation to the ODE solution above (see Remark 4.8 for further discussion). However, the authors in [15, 20] did not prove the solution’s optimality as stated in Corollary 3.14. Another similar approach can be found in Section III in [4], where an approximate parallel transport scheme was used to construct $u_{k}$. Different matching conditions are imposed to obtain assignments of $u_{k}$ that are only continuous in $k$, producing Wannier functions that decay as $1/x^{2}$, which are then used as inputs for the general local optimization approach in [15].
@@ 4.1.1 Computing the Fourier interpolant of the potential
We then evaluate $V(t_{j})$ for all $j=1,2,\ldots,2M+1$. The $12M+$ 1 Fourier coefficients $\{\widehat{V}_{j}\}_{j=-M}^{M}$ in (53) are then computed via a discrete Fourier transform. We then form the potential matrix $\mathbf{V}\in\mathbb{C}^{(2M+1)\times(2M+1)}$ with entries given by
@@ 4.1.2 Obtaining an initial condition
where $E_{0}$ is the $\ell$th smallest eigenvalue and $\mathbf{y}^{(0)}$ is the corresponding eigenvector, normalized as in (55) such that
@@
This can be done using the standard QR algorithm (cf. e.g., [11, Section 8.3]).
@@
Remark 4.4.
If the standard QR algorithm is used, this procedure takes $O(M^{3})$ floating point operations. It should be noted that the coefficient matrix in (111) is highly structured. In particular, it consists of a diagonal matrix plus a Toeplitz matrix, which can likely be exploited to accelerate this step.
@@ 4.1.3 Solving the initial value problem
where $E_{0}$ and $\mathbf{y}^{(0)}$ were obtained in (111). We would like to solve this initial value problem from $-\Omega/2$ to $\Omega/2$. We note that by (112) $\|\mathbf{y}(k)\|=1/a$ for $k\in[-\Omega/2,\Omega/2]$.
@@
must be inverted. In this paper, we compute $\mathbf{\Theta}(k)^{\dagger}$ by first computing a singular value decomposition of $\mathbf{\Theta}(k)$, discarding the mode corresponding to the smallest singular value (see Remark 2.9), and applying the inverse of the decomposition directly to the right-hand side (cf. e.g., [1, Section 2.6.1]). It is not as efficient as the approach in Remark 2.10, but this is not significant for one dimension problems.
Remark 4.5.
In our numerical experiments, we use the standard 4th-order Runge–Kutta method to solve the initial value problem for simplicity (cf. e.g., [18, Section 12.5]); the resulting solution produces a solution that converges at rate $O(K^{-4})$. Even higher order, for example $O(K^{-12})$, can be achieved by switching to a spectral deferred scheme (cf. [10]).
Remark 4.6.
At each time-step, the dominant cost is that of computing the singular value decomposition of $\mathbf{\Theta}(k)$, which consists of $O(M^{3})$ floating point operations. Since $K$ time steps are taken, this stage of our procedure requires $O(KM^{3})$ floating point operations.
Remark 4.7.
The explicit dependence of (114) on $E(k)$ and the ODE for the energy (115) can be eliminated altogether by replacing $E(k)$ with the Rayleigh quotient
@@ 4.1.4 Correcting the phase
According to Lemma 3.10, we can determine $\phi_{\rm zak}$ from
$\mathbf{y}^{(0)}_{m+1}=e^{\mathrm{i}\phi_{\rm zak}}\mathbf{y}^{(K)}_{m},$ (122)
for any $m\in\{1,2,\ldots,2M\}$ to fix the jumps. Moreover, the condition in (100) allows us to determine $\phi_{0}$ by
$e^{-\mathrm{i}\phi_{0}}\mathbf{y}^{(0)}_{m}=e^{\mathrm{i}\phi_{0}}\overline{\mathbf{y}}^{(K)}_{2M+2-m}$ (123)
@@
$\phi(k_{j})=\frac{\phi_{\rm zak}}{\Omega}k_{j},\quad j=1,2,\ldots,K.$ (124)
@@
$\widetilde{\mathbf{y}}^{(j)}=e^{-\mathrm{i}\phi_{0}}e^{-\mathrm{i}\phi(k_{j})}\mathbf{y}^{(j)},\quad j=1,2,\ldots,K,$ (125)
@@ 4.1.6 Evaluate the Wannier function
Remark 4.9.
Since the integrand in (48) is analytic and exponentially decaying as $\xi\to\pm\infty$, the error in the trapezoidal rule approximation decays with respect to $K$ as $O(e^{-cK})$ for some $c>0$ (cf. e.g., [19]).
Remark 4.10.
The cost of direct evaluation of (128) for each $x$ is $O(KM)$. The evaluation of (128) for a collection of points can be accelerated using the fast Fourier transform or non-uniform fast Fourier transform (cf. e.g., [12]).
@@ 5 Detailed description of the procedure
    Solve the eigenvalue problem (111) for the $\ell$th smallest eigenvalue $E^{(0)}$ and corresponding eigenvector $\mathbf{y}^{(0)}$ using the QR algorithm and ensuring $\|\mathbf{y}^{(0)}\|=1$.
@@ 3. End Do
      Set $\widetilde{\mathbf{y}}^{(j)}=\exp(-\mathrm{i}\phi_{0})\exp\left(-\mathrm{i}\phi_{\rm zak}k_{j}/\Omega\right)\mathbf{y}^{(j)}$.
@@ 6.1 Gaussian potential
Figure 1: Plot of potential (left) and $E^{(j)}(k)$ (right) for $j=1,2,3$ (red, blue, green, resp.)
@@ 6.2 Asymmetric potential
$V(x)=\frac{1}{4}\left(1+2\sin(2\Omega x)+3e^{\cos(\Omega x)}\right).$ (134)
@@ 7 Conclusion and generalization
Since Kato’s theorem (Theorem 2.3) guarantees analyticity even when eigenvalues become degenerate in one dimension, the approach in this paper has been easily extended to isolated multiband cases in one dimension. Similarly, globally optimal solutions can be computed analytically. For higher dimensions, topological phenomena show up in the construction of Wannier functions and new ideas beyond this paper are required. Nonetheless, the algorithms in one dimension are the basic building block for higher-dimensional cases. Such extensions to higher dimensions have been worked out and are now in preparation for publication.
@@ Acknowledgements
The authors thank John Schotland for drawing their attention to this subject. The authors also thank Vladimir Rokhlin for helpful discussions pertaining to this work.

### v4

@@ 4 Constructing optimally localized Wannier functions
The construction can be divided into three stages. First, we compute the Fourier series of $V$ as in (53) and transform (32) into its Fourier version in (56). Then, we compute the energy $E_{0}$ and the eigenvector $\mathbf{y}_{0}$ at $k=-\Omega/2$ for the selected band. Second, we solve the ODE defined by (60) and (61) in $\Omega/2,\Omega/2]$ with initial conditions given by $E_{0}$ and $\mathbf{y}_{0}$ at $k=-\Omega/2$. From $\mathbf{y}(-\Omega/2)$ and $\mathbf{y}(\Omega/2)$, we determine a constant $\varphi_{0}$ to enforce (100) and the Zak phase $\varphi_{\rm zak}$ in (68). This is followed by carrying out a gauge transform to obtain the new $\widetilde{\mathbf{y}}$ in (76) with the Berry phase $\varphi$ given by (80). Using the definition of $\widetilde{\mathbf{y}}$ in (77), we obtain a function $\widetilde{\alpha}$ in $\mathbb{R}$ that defines the Fourier transform of the Wannier function by (49). Finally, we Fourier transform $\widetilde{\alpha}$ to obtain the Wannier function $W_{0}$ by (48). This Wannier function is guaranteed to be optimal by Theorem 3.10 and Corollary 3.12, and real by Theorem 3.13.
@@ 4.1.1 Computing the Fourier interpolant of the potential
We then evaluate $V(t_{j})$ for all $j=1,2,\ldots,2M+1$. The $2M+1$ Fourier coefficients $\{\widehat{V}_{j}\}_{j=-M}^{M}$ in (53) are then computed via a discrete Fourier transform. We then form the potential matrix $\mathbf{V}\in\mathbb{C}^{(2M+1)\times(2M+1)}$ with entries given by
@@ 4.1.2 Obtaining an initial condition
[…]
@@
This can be done using the standard QR algorithm (cf. e.g., [11, Section 8.3]). This procedure would then take $O(M^{3})$ floating point operations. However, it should be noted that the coefficient matrix in (111) is highly structured. In particular, it consists of a diagonal matrix plus a Toeplitz matrix, which can likely be exploited to accelerate this step.
@@ 4.1.3 Solving the initial value problem
where $E_{0}$ and $\mathbf{y}^{(0)}$ were obtained in (111). We would like to solve this initial value problem from $\Omega/2$ to $\Omega/2$. We note that by (112) $\|\mathbf{y}(k)\|=1/a$ for $k\in[-\Omega/2,\Omega/2]$.
@@
must be inverted. In this paper, we compute $\mathbf{\Theta}(k)^{\dagger}$ by first computing a singular value decomposition of $\mathbf{\Theta}(k)$, discarding the mode corresponding to the smallest singular value, and applying the inverse of the decomposition directly to the right-hand side (cf. e.g., [1, Section 2.6.1]). It is not as efficient as the approach in Remark 2.8, but this is not significant for one-dimensional problems.
Since this problem is not stiff, any time-stepping scheme may be chosen. In our numerical experiments, we use the standard 4th-order Runge–Kutta method to solve the initial value problem for simplicity (cf. e.g., [20, Section 12.5]); the resulting solution produces a solution that converges at rate $O(K^{-4})$. Even higher order, for example $O(K^{-12})$, can be achieved by switching to a spectral deferred scheme (cf. [10]). At each time-step, the dominant cost is that of computing the singular value decomposition of $\mathbf{\Theta}(k)$, which consists of $O(M^{3})$ floating point operations. Since $K$ time steps are taken, this stage of our procedure requires $O(KM^{3})$ floating point operations.
We note that the explicit dependence of (114) on $E(k)$ and the ODE for the energy (115) can be eliminated altogether by replacing $E(k)$ with the Rayleigh quotient
@@ 4.1.4 Correcting the phase
According to Lemma 3.9, we can determine $\varphi_{\rm zak}$ from
$\mathbf{y}^{(K)}_{m}=e^{\mathrm{i}\varphi_{\rm zak}}\mathbf{y}^{(0)}_{m+1}$, (122)
for any $m\in\{1,2,\ldots,2M\}$ to fix the jumps. Moreover, the condition in (100) allows us to determine $\varphi_{0}$ by
$e^{-\mathrm{i}\varphi_{0}}\mathbf{y}^{(0)}_{m}=e^{\mathrm{i}\varphi_{0}}\overline{\mathbf{y}}^{(K)}_{2M+2-m}$ (123)
@@
$\varphi(k_{j})=\frac{\varphi_{\rm zak}}{\Omega}k_{j},\quad j=1,2,\ldots,K.$ (124)
@@
$\widetilde{\mathbf{y}}^{(j)}=e^{-\mathrm{i}\varphi_{0}}e^{-\mathrm{i}\varphi(k_{j})}\mathbf{y}^{(j)},\quad j=1,2,\ldots,K,$ (125)
@@ 4.1.6 Evaluate the Wannier function
The rate of convergence of the sum in (128) with respect to $K$ and $M$ depends on the error in the trapezoidal rule approximation to the integral in (48). The integrand in (48) is analytic and decays as $\xi\to\pm\infty$. Thus, it follows from the Euler–Maclaurin formula that the precise rate of convergence will depend on the rate at which the integrand goes to zero, which in turn is determined by the smoothness of the potential (see Remark 4.2). When the integrand decays exponentially $\xi\to\pm\infty$, the error in the trapezoidal rule approximation decays with respect to $K$ as $O(e^{-cK})$ for some $c>0$ respect to $K$ as $O(e^{-cK})$ for some $c>0$ (cf. e.g., [21]).
We note that while the cost of direct evaluation of (128) for each $x$ is $O(KM)$, the evaluation of (128) for a collection of points can be accelerated using the fast Fourier transform or non-uniform fast Fourier transform (cf. e.g., [12]).
@@ 5 Detailed description of the procedure
[…]
@@ 3. End Do
      Set $\widetilde{\mathbf{y}}^{(j)}=\exp(-\mathrm{i}\varphi_{0})\exp\left(-\mathrm{i}\varphi_{\rm zak}k_{j}/\Omega\right)\mathbf{y}^{(j)}$.
@@ 6.1 Gaussian potential
Figure 1: Plot of potential (left) and $E^{(j)}(k)$ (right) for $j=1,2,3$ (bottom, middle, top, resp.)
@@
Figure 2: $W_{0}^{(j)}(x)$ (left) and $\log_{10}(|W_{0}^{(j)}(x)|)$ (right) for $j=1,2,3$ (top, middle, bottom, resp.) We note that while all the Wannier functions decay exponentially, the constants in this rate depend strongly on the band, as demonstrated by the different horizontal axes scales.
@@ 6.2 Asymmetric potential
Figure 3: Plot of potential (left) and $E^{(j)}(k)$ (right) for $j=1,2,3$ (bottom, middle, top, resp.)
@@
$V(x)=-\frac{1}{4}\left(1+2\sin(2\Omega x)+3e^{\cos(\Omega x)}\right).$ (134)
@@
Again, the different horizontal axes scales show that the constants in the decay depend on the band.
@@ 7 Conclusion and generalization
The key challenge in extending the scheme presented here to the multiband case is that, instead of scalar phases as in the single-band case in this paper, all possible gauge choices form a family of unitary matrices, i.e. the matrix $U(N_{b})$ group where $N_{b}$ is the number of bands considered. While this makes the analysis somewhat more involved, the underlying concepts remain largely the same, and globally optimal solutions can still be computed analytically. In higher dimensions, topological phenomena emerge in the construction of Wannier functions, requiring ideas beyond those presented in this paper, as demonstrated in [24] for matrix models. Nonetheless, the one-dimensional algorithms serve as the fundamental building blocks for higher-dimensional cases. Extensions to higher dimensions have been worked out and are currently being prepared for publication.
@@ Acknowledgements
The authors are grateful to John Schotland for drawing their attention to this subject, and to Vladimir Rokhlin for valuable discussions pertaining to this work. The authors also thank the anonymous referees for their helpful suggestions.